# On the Hardness of Approximating Distributions with Probabilistic Circuits

### v1

@@ Abstract
A fundamental challenge in probabilistic modeling is balancing expressivity and tractable inference. Probabilistic circuits (PCs) aim to directly address this tradeoff by imposing structural constraints that guarantee efficient inference of certain queries while maintaining expressivity. Since inference complexity on PCs depends on circuit size, understanding the size bounds across circuit families is key to characterizing the tradeoff between tractability and expressive efficiency. However, expressive efficiency is often studied through exact representations, where exactly encoding distributions while enforcing various structural properties often incurs exponential size blow-ups. Thus, we pose the following question: can we avoid such size blow-ups by allowing some small approximation error? We first show that approximating an arbitrary distribution with bounded $f$-divergence is $\mathsf{NP}$-hard for any model that can tractably compute marginals. We then prove an exponential size gap for approximation between the class of decomposable PCs and additionally deterministic PCs.
@@ 1 Introduction
Generative models have shown remarkable success in capturing complex distributions [21, 24, 28, 22, 38]; yet, despite their expressivity, they often do not support efficient computation of fundamental probabilistic queries, which are critical for inference in domains such as healthcare [33], neuro-symbolic AI [36], algorithmic fairness [7], and environmental science [3]. Probabilistic circuits address this by balancing expressivity and tractable inference. Naturally, there have been many works characterizing the expressive efficiency of different circuit classes [16, 1, 6, 40, 41, 17]. However, how structural constraints affect the ability to approximate distributions remains underexplored.
@@
This paper studies the following fundamental question: does allowing small approximation error alleviate the exponential separation between circuit classes observed in exact modeling, or does hardness persist even in the approximate setting? Our motivation is two-fold. (1) Showing that certain distributions cannot be approximated within a bounded distance compactly by a family of PCs implies that any learning algorithm whose hypothesis space is that family of PCs would fail to learn the distribution with a bounded approximation error. (2) Moreover, PCs can also be used to perform inference on other probabilistic models (such as Bayesian networks or probabilistic programs) by compiling them into PCs then performing inference on the compiled circuits [2, 23, 14, 5, 20]. This suggests the following approximate inference scheme: approximately compile a probabilistic model into a PC then run efficient exact inference on the approximately compiled PC. Thus, if we can bound the distance between the target distribution and approximate model, we can hope to provide guarantees on the approximate inference results as well.
@@
Our main contributions are as follows: (1) we prove that it is $\mathsf{NP}$-hard to approximate distributions within a bounded $f$-divergence using any model that tractably computes marginals, via a reduction from SAT (Theorem 3.3 and 3.4); (2) we derive an unconditional, exponential separation between decomposable PCs and decomposable & deterministic PCs for approximate modeling (Theorem 4.1); (3) we study the relationship between bounds on divergence measures for approximate modeling and approximation errors for marginal and maximum-a-posteriori (MAP) inference, characterizing when one is or is not sufficient to guarantee the other (Section 3 and further discussion in Appendix B).
@@ 2 Preliminaries
Probabilistic Circuits
A probabilistic circuit (PC) recursively encodes a joint probability distribution over random variables through a parameterized directed acyclic graph (DAG), composed of 3 types of nodes: leaf, product $\otimes$, and sum nodes $\oplus$. PCs are closely related to logical circuits in the knowledge compilation literature [16]. Logical circuits encode Boolean functions as directed acyclic graphs (DAGs) consisting of AND ($\wedge$) and OR ($\vee$) gates—analogous to $\otimes$ and $\oplus$ nodes, respectively—with positive and negative literals as leaf nodes. We can characterize different families of both probabilistic and logical circuits based on their structural properties, such as decomposability and determinism. These circuit properties allow tractable (polytime) computation of various queries); for instance, decomposable PCs admit linear-time marginal inference while deterministic and decomposable PCs additionally allow linear-time maximum-a-posteriori (MAP) inference. There is a rich literature studying different families of circuit representations in terms of their tractability for inference and operations, as well as their relative succinctness (expressive efficiency) for both exact and approximate compilation, which we will leverage for our hardness results and size lower bounds on PCs. See Appendix A for further background on PCs.
Measures of Difference between Probability Distributions
To study the hardness of approximating probability distributions, we must first measure how “good” an approximation is. We focus on the class of $f$-divergences.
@@
Commonly used $f$-divergences include the Kullback-Leibler divergence, $\chi^{2}$-divergence, and the total variation distance, which is especially relevant to our results.
@@
To more compactly describe PCs that approximate distributions under some bounded distance, we use the following.
@@
Let $P,Q$ be some probability measures and $D$ be some distance we want to minimize between $P$ and $Q$. We say that $Q$ is an $\epsilon$-$D$-Approximator of $P$ if $D(P||Q)<\epsilon$ for some $\epsilon>0$.
We refer to any circuit $Q$ such that it approximates our target distribution $P$ within $D_{\mathsf{TV}}(P\|Q)<\epsilon$ a $\epsilon$-$\mathcal{D}_{\mathsf{TV}}$-Approximator. The majority of our results are derived using properties of the total variation distance, due to its properties as a distance metric. To extend our results to $f$-divergences, we utilize the class of $k$-convex f-divergences, which provides an upper bound on the total variation distance.
We can define a $k$-convex f-divergence [27] as a ${}\mathbb{R}\cup\{\infty\}$-valued function $f$ on a convex set $K\subseteq\mathbb{R}$ is $k$-convex when $x,y\in K$ and $t\in[0,1]$ implies $f((1-t)x+ty)\leq(1-t)f(x)+tf(y)-kt(1-t)\frac{(x-y)^{2}}{2}.$ When $f$ is twice differentiable, this is equivalent to $f^{\prime\prime}(x)\geq k$ for $x\in K$. In the case that $k=0$ this reduces to the normal notion of convexity. An $f$-divergence $D_{f}$ is $k$-convex over an interval $K$ for $k\geq 0$ when the function $f$ is $k$-convex on $K$.
We provide a table in Appendix C summarizing which $f$-divergence measures are $k$-convex and for which value of $k$. Throughout this paper, we express approximation bounds using $k$-convex $f$-divergences as they naturally encapsulate bounds on many common distance measures. For instance, for any $k$-convex $f$-divergence between $P$ and $Q$, we have that $D_{\mathsf{TV}}(P\|Q)^{2}<{D_{f}(P||Q)}/{k}$ [27]. For KL-divergence, which is the most commonly used objective for learning probabilistic models, we can use a form of Pinsker’s inequality [37] to obtain $D_{\mathsf{TV}}(P\|Q)<\sqrt{\frac{1}{2}D_{\mathsf{KL}}(P\|Q)}$.
@@ 3 Approximate Modeling with Tractable Marginals is $\mathsf{NP}$-hard
Most works characterizing the expressive efficiency of different circuit classes have been concerned with exact representations [16, 1, 6, 40, 41, 17]. While Chubarian and Turán [10] and De Colnet and Mengel [18] have recently studied the ability (and hardness) of logical circuit classes to compactly approximate Boolean functions, to the best of our knowledge, our results are the first to show hardness of compactly approximating probability distributions for different families of tractable PCs.
As noted earlier, the complexity of approximately modeling distributions with PCs is valuable for understanding: (1) potential limitations in the hypothesis space of PC learning algorithms, and (2) the feasibility of approximate inference with guarantees through approximate compilation. This section aims to answer this, focusing on probabilistic models that are tractable for marginal queries. We first show that a form of approximate marginal inference using this scheme requires a non-trivial bound on the total variation distance between the target distribution and the approximate model, and prove that finding such an approximator is $\mathsf{NP}$-hard.
We consider relative (multiplicative) approximation of marginal queries. Let $P(\mathbf{X})$ be a probability distribution over a set of variables $\mathbf{X}$. Then we say another distribution $Q(\mathbf{X})$ is a relative approximator of marginals of $P$ w.r.t. $0\leq\epsilon\leq 1$ if: $\frac{1}{1+\epsilon}\leq\frac{P(\mathbf{y})}{Q(\mathbf{y})}\leq 1+\epsilon$ for every assignment $\mathbf{y}$ to a subset $Y\subseteq\mathbf{X}$. Relative approximation of marginals is known to be $\mathsf{NP}$-hard for Bayesian networks [12], thus it immediately follows that approximately representing arbitrary distributions using polynomial-sized PCs tractable for marginals such that the PC is a relative approximator of all marginals is also $\mathsf{NP}$-hard. However, approximating all marginal queries is quite a strong condition, and we may still want to closely approximate distributions as they could be useful in approximating some marginal queries.
We show that relative approximation for all marginal queries implies a non-trivially bounded total variation distance.
@@
As $Q$ is a relative approximator of $P$, for all assignment $\mathbf{x}$ we have $\frac{1}{1+\epsilon}\leq\frac{P(\mathbf{x})}{Q(\mathbf{x})}\leq 1+\epsilon$ which implies $\left\lvert P(\mathbf{x})-Q(\mathbf{x})\right\rvert\leq\epsilon\min(P(\mathbf{x}),Q(\mathbf{x}))$. Therefore, $D_{\mathsf{TV}}(P\|Q)=\frac{1}{2}\sum_{\mathbf{x}}\left\lvert P(\mathbf{x})-Q(\mathbf{x})\right\rvert\leq\frac{1}{2}\sum_{\mathbf{x}}\epsilon\min(P(\mathbf{x}),Q(\mathbf{x}))\leq\frac{\epsilon}{2}$.
In other words, $D_{\mathsf{TV}}(P\|Q)\leq\epsilon/2$ is a necessary condition for $Q$ to be a relative approximator of marginals of $P$ w.r.t. $\epsilon$. However, this is not a sufficient condition as shown below.
@@
We prove the above proposition by explicitly constructing a family of distributions $\mathcal{Q}$ such that every $Q\in\mathcal{Q}$ is an $\epsilon$-$D_{f}$-approximator for any arbitrary $\epsilon>0$ and distribution $P$ yet $P(\mathbf{x})/Q(\mathbf{x})$ can be arbitrarily large for some $\mathbf{x}$. See Appendix E.1 for the full construction.
Since approximating a distribution $P$ with a bounded TV distance is a necessary but insufficient condition for the $\mathsf{NP}$-hard problem of relative approximation of marginals, this raises the question whether it is still possible to efficiently approximate the distribution $P$ with a compact PC $Q$ that is tractable for marginals. We next answer this in the negative.

### v2

# On the Hardness of Approximating Distributions with Tractable Probabilistic Models

@@ Abstract
A fundamental challenge in probabilistic modeling is to balance expressivity and inference efficiency. Tractable probabilistic models (TPMs) aim to directly address this tradeoff by imposing constraints that guarantee efficient inference of certain queries while maintaining expressivity. In particular, probabilistic circuits (PCs) provide a unifying framework for many TPMs, by characterizing families of models as circuits satisfying different structural properties. Because the complexity of inference on PCs is a function of the circuit size, understanding the size requirements of different families of PCs is fundamental in mapping the trade-off between tractability and expressive efficiency. However, the study of expressive efficiency of circuits are often concerned with exact representations, which may not align with model learning, where we look to approximate the underlying data distribution closely by some distance measure. Moreover, due to hardness of inference tasks, exactly representing distributions while supporting tractable inference often incurs exponential size blow-ups. In this paper, we consider a natural, yet so far underexplored, question: can we avoid such size blow-up by allowing for some small approximation error? We study approximating distributions with probabilistic circuits with guarantees based on $f$-divergences, and analyze which inference queries remain well-approximated under this framework. We show that approximating an arbitrary distribution with bounded $f$-divergence is $\mathsf{NP}$-hard for any model that can tractably compute marginals. In addition, we prove an exponential size gap for approximation between the class of decomposable PCs and that of decomposable and deterministic PCs.
@@ 1 Introduction
The expressive power of probabilistic and generative models has increased rapidly in recent years: from Bayesian networks [18], GANs [26], VAEs [29], and normalizing flows [35], to diffusion models [27] and transformers [46] have demonstrated remarkable success in capturing complex distributions. Yet, despite their expressivity, many of these models do not support efficient computation of fundamental probabilistic queries, such as marginals and conditionals, which are critical for inference in domains such as healthcare [40], neuro-symbolic AI [43], environmental science [5], and algorithmic fairness [9].
Tractable probabilistic models—such as probabilistic circuits [8], probabilistic generating circuits [50], determinantal point processes [1], and more—address the need for probabilistic queries by balancing expressivity and tractable inference, achieved through enforcing constraints on the models. To more compactly describe the constraints enforced, we focus on probabilistic circuits, which also have extensive literature on the conditions of tractability [8]. While these structural constraints enable efficient inference, they introduce a tradeoff by potentially affecting the models’ ability to compactly represent distributions (i.e., their expressive power). Naturally, there have been many works characterizing the expressive efficiency of different circuit classes [19, 2, 8, 48, 49, 20]—i.e., their ability to compactly and exactly represent certain classes of functions or distributions. However, the approximate case—how structural constraints affect the ability to approximately represent distributions, remains comparatively underexplored. This motivates a shift in focus from exact to approximate modeling: representing a distribution approximately within a small distance under some metric.
@@
This shift raises a fundamental question: does allowing a small approximation error alleviate the exponential separation between circuit classes observed in exact modeling, or does hardness persist even in the approximate setting? Our motivation for this study is two-fold. (1) In learning probabilistic circuits from finite data, often the goal is not necessarily to exactly represent some known distribution but rather to approximate it as closely as possible with a PC of reasonable size. Thus, showing that certain distributions cannot be approximated within a bounded distance by a compact PC satisfying some structural properties implies that any learning algorithm whose hypothesis space is that family of PCs would fail to learn the distribution with a bounded approximation error. (2) Moreover, probabilistic circuits can also be used to perform inference on other probabilistic models (such as Bayesian networks or probabilistic programs) by compiling them into PCs then running efficient inference on the compiled circuits [4, 28, 16, 7, 25]. This suggests the following approximate inference scheme: approximately compile a probabilistic model into a PC then run efficient exact inference on the approximately compiled PC. Moreover, if we could bound the distance between the target distribution and approximate model, we can hope to provide guarantees on the approximate inference results as well.
@@
Our main contributions are as follows: (1) we prove that it is $\mathsf{NP}$-hard to approximate distributions within a bounded $f$-divergence using any model that supports tractable marginals, with proof via a reduction from SAT (Theorem 3.4 and 3.5); (2) we derive an unconditional, exponential separation between decomposable PCs and decomposable & deterministic PCs for approximate modeling (Theorem 4.1); (3) we study the relationship between bounds on divergence measures for approximate modeling and approximation errors for marginal and maximum-a-posteriori (MAP) inference, characterizing when one is or is not sufficient to guarantee the other (Sections 3.1 and 5).
@@ 2 Preliminaries
We use uppercase letters ($X$) to denote random variables and lowercase letters ($x$) to denote assignments to these random variables. Sets of random variables and assignments are denoted using bold letters ($\mathbf{X}\text{ and }\mathbf{x}$). The accepting models (i.e., satisfying assignments) of a Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ over $n$ variables is denoted by $f^{-1}(1)$. The number of accepting models of $f$ is referred to as $\operatorname{MC}(f)$, a shorthand for its model count. Moreover, a Boolean function $f$ is the support of a distribution $P$, if $P$ is non-zero only over the models of $f$.
2.1 Probabilistic Circuits
Definition 2.1 (Probabilistic circuits).
A probabilistic circuit (PC) $\mathcal{C}:=(\mathcal{G},\theta)$ represents a joint probability distribution $p(\mathbf{X})$ over random variables $\mathbf{X}$ through a directed acyclic graph (DAG) $\mathcal{G}$ parameterized by $\theta$. The DAG is composed of 3 types of nodes: leaf, product $\otimes$, and sum $\oplus$ nodes. Every leaf node in $\mathcal{G}$ is an input, and every internal node receives inputs from its children $\operatorname{in}(n)$. The scope of a given node, $\phi(n)$, is a recursively defined function which associates to each unit $n$ a subset of $\mathbf{X}$: for each non-input unit $n$, $\phi(n)=\cup_{c\in\operatorname{in}(n)}\phi(c)$, and the scope of a leaf node is a single variable in $\mathbf{X}$. Naturally, the scope of the root node is $\mathbf{X}$. Each node $n$ of a PC is then recursively defined as:
@@ 2.1 Probabilistic Circuits
where $\theta_{n,c}\in[0,1]$ is the parameter associated with the edge connecting nodes $n,c$ in $\mathcal{G}$, and $\sum_{c\in\operatorname{in}(n)}\theta_{n,c}=1.$ In this paper, we assume $l(x)$ at a leaf node is a Boolean indicator function: i.e., $\mathbbm{1}[x=1]$ or $\mathbbm{1}[x=0]$. The distribution represented by the circuit is the output at its root node.
A key characteristic of PCs is that imposing certain structural properties on the circuit enables tractable (polytime) computation of various queries. In this paper we focus on two families of PCs: those that are tractable for marginal inference and for maximum-a-posteriori (MAP) inference.
The class of marginal queries of a joint distribution $p(\mathbf{X})$ over variables $\mathbf{X}$ refers the set of functions that can compute $p(\mathbf{y})$ for some assignment $\mathbf{y}$ for $\mathbf{Y}\subseteq\mathbf{X}$. Marginalization is a fundamental statistical operation which enables reasoning about subsets of variables, essential for tasks such as decision making, learning, and predicting under uncertainty. While marginal inference is $\mathsf{\#P}$-hard in general [39, 13], the family of PCs satisfying the following structural conditions admit tractable marginal inference—specifically in linear time in the size of the circuit [17].
Definition 2.2 (Smoothness and decomposability).
A sum unit is smooth if its children have identical scopes: $\phi(c)=\phi(n),\,\forall c\in\operatorname{in}(n)$. A product unit is decomposable if its children have disjoint scopes: $\phi(c_{i})\cap\phi(c_{j})=\emptyset,\,\forall c_{i}\neq c_{j}\in\operatorname{in}(n)$. A PC is smooth and decomposable iff every sum unit is smooth and every product unit is decomposable.
[…]
Definition 2.3 (Determinism).
A sum node is deterministic if, for any fully-instantiated input, the output of at most one of its children is nonzero. In other words, the supports of its children are mutually disjoint. A PC is deterministic iff all of its sum nodes are deterministic.
Logical circuits
Probabilistic circuits are closely related to logical circuits in the knowledge compilation literature [19]. Logical circuits encode Boolean functions as directed acyclic graphs consisting of AND ($\wedge$) and OR ($\vee$) gates with positive and negative literals as leaf nodes. We can also characterize different families of logical circuits based on their structural properties: e.g., decomposable negation normal forms (DNNFs) and deterministic decomposable negation normal forms (d-DNNFs). [^3] Structural conditions are same as before (Definitions 2.2 and 2.3), except that smoothness and determinism apply to OR gates and decomposability to AND gates. There is a rich literature studying different logical circuit families in terms of their tractability for inference and operations, as well as their relative succinctness (expressive efficiency) for both exact and approximate compilation, which we will leverage for our hardness results and size lower bounds on probabilistic circuits.
2.2 Measures of Difference between Probability Distributions
To study the hardness of approximating probability distributions, we first need to be able to measure how “good” an approximation is. In particular, we focus on the class of $f$-divergences.
@@ 2.2 Measures of Difference between Probability Distributions
Commonly used $f$-divergences include the Kullback-Leibler divergence, $\chi^{2}$-divergence, and total variation distance. The total variation distance is especially relevant to our results.
@@
We introduce the following notion to describe probabilistic models that approximate distributions within some bounded distance.
@@
Let $P,Q$ be two probability distributions and $D$ be a distance measure between distributions. We say that $Q$ is an $\epsilon$-$D$-approximator of $P$ if $D(P||Q)<\epsilon$ for some $\epsilon>0$.
For instance, we refer to a probabilistic circuit $Q$ that approximates our target distribution $P$ such that $D_{\mathsf{TV}}(P\|Q)<\epsilon$ a $\epsilon$-$\mathcal{D}_{\mathsf{TV}}$-approximator. The majority of our results are derived using properties of the total variation distance, due to its nice properties as a distance metric. To extend our results to other $f$-divergences, we utilize the following class, which provides an upper bound on the total variation distance.
Definition 2.7 ($k$-convex $f$-divergence [34]).
A $\mathbb{R}\cup\{\infty\}$-valued function $f$ on a convex set $K\subseteq\mathbb{R}$ is $k$-convex if $x,y\in K$ and $t\in[0,1]$ implies $f((1-t)x+ty)\leq(1-t)f(x)+tf(y)-kt(1-t)\frac{(x-y)^{2}}{2}.$ When $f$ is twice differentiable, this is equivalent to $f^{\prime\prime}(x)\geq k$ for all $x\in K$. In the case that $k=0$ this reduces to the normal notion of convexity. An $f$-divergence $D_{f}$ is $k$-convex over an interval $K$ for $k\geq 0$ if the function $f$ is $k$-convex on $K$.
We provide a table in Appendix A.1 summarizing which $f$-divergence measures are $k$-convex and for which value of $k$. Throughout this paper, we express approximation bounds using $k$-convex $f$-divergences as they naturally encapsulate bounds on many common distance measures. For instance, for any $k$-convex $f$-divergence between $P$ and $Q$, we have that $D_{\mathsf{TV}}(P\|Q)^{2}<{D_{f}(P||Q)}/{k}$ [34]. As we will see later, the bounds on the TV distance can naturally be connected to guarantees for approximation inference. For KL-divergence, which is the most commonly used objective for learning probabilistic models, we can use Pinsker’s inequality [45] to obtain $D_{\mathsf{TV}}(P\|Q)<\sqrt{\frac{1}{2}D_{\mathsf{KL}}(P\|Q)}$.
@@ 3 Approximate Modeling with Tractable Marginals is $\mathsf{NP}$-hard
Most works characterizing the expressive efficiency of different circuit classes have been concerned with exact representations [19, 2, 8, 48, 49, 20]. While Chubarian and Turán [12] and De Colnet and Mengel [21] have recently studied the ability (and hardness) of logical circuit classes to compactly approximate Boolean functions, to the best of our knowledge, our results are the first to show hardness of compactly approximating probability distributions using different families of tractable PCs.
As discussed previously, the complexity of approximately modeling distributions with PCs is valuable for understanding: (1) potential limitations in the hypothesis space of PC learning algorithms, and (2) the feasibility of approximate inference with guarantees through approximate compilation. This section aims to answer this focusing on probabilistic models that are tractable for marginal queries. We first show that a form of approximate marginal inference using this scheme requires a non-trivial bound on the total variation distance between the target distribution and the approximate model, and next prove that finding such an approximator is $\mathsf{NP}$-hard.
3.1 Relative Approximation of Marginals
We consider relative approximation [^4] Also called multiplicative approximation or approximation within a factor. of marginal queries. Let $P(\mathbf{X})$ be a probability distribution over a set of variables $\mathbf{X}$. Then we say another distribution $Q(\mathbf{X})$ is a relative approximator of marginals of $P$ w.r.t. $0\leq\epsilon\leq 1$ if: $\frac{1}{1+\epsilon}\leq\frac{P(\mathbf{y})}{Q(\mathbf{y})}\leq 1+\epsilon$ for every assignment $\mathbf{y}$ to subset of variables $Y\subseteq\mathbf{X}$. Relative approximation is often considered for approximate inference of graphical models and the closely related approximate (weighted) model counting [23, 3]. We first show that relative approximation for all marginal queries implies a non-trivially bounded total variation distance.
@@ 3.1 Relative Approximation of Marginals
Proof.
As $Q$ is a relative approximator of $P$, for all assignment $\mathbf{x}$ we have that $\frac{1}{1+\epsilon}\leq\frac{P(\mathbf{x})}{Q(\mathbf{x})}\leq 1+\epsilon$ which implies $\left\lvert P(\mathbf{x})-Q(\mathbf{x})\right\rvert\leq\epsilon\min(P(\mathbf{x}),Q(\mathbf{x}))$. Therefore, $D_{\mathsf{TV}}(P\|Q)=\frac{1}{2}\sum_{\mathbf{x}}\left\lvert P(\mathbf{x})-Q(\mathbf{x})\right\rvert\leq\frac{1}{2}\sum_{\mathbf{x}}\epsilon\min(P(\mathbf{x}),Q(\mathbf{x}))\leq\frac{\epsilon}{2}$. ∎
In other words, $D_{\mathsf{TV}}(P\|Q)\leq\epsilon/2$ is a necessary condition for $Q$ to be a relative approximator of marginals of $P$ w.r.t. $\epsilon$. However, it is still not a sufficient condition as shown below.
@@
We prove the above proposition by explicitly constructing a family of distributions $\mathcal{Q}$ such that every $Q\in\mathcal{Q}$ is an $\epsilon$-$D_{f}$-approximator for any arbitrary $\epsilon>0$ and distribution $P$ yet $P(\mathbf{x})/Q(\mathbf{x})$ can be arbitrarily large for some $\mathbf{x}$. See Appendix A.3.1 for the full construction.
It is known that relative approximation of marginals is $\mathsf{NP}$-hard for Bayesian networks [14]. [^5] In fact, there also exists no randomized polynomial time algorithm for relative approximation of marginals unless $\mathsf{RP}=\mathsf{NP}$ [14]. Thus, it immediately follows that approximately representing arbitrary distributions using polynomial-sized PCs tractable for marginals (e.g., decomposable PCs) such that the PC is a relative approximator of all marginals is also $\mathsf{NP}$-hard. However, approximating all marginal queries is quite a strong condition, and we may still want to closely approximate distributions as they could be useful in approximating some marginal queries. In particular, because approximating a distribution $P$ with a bounded TV distance is a necessary but not sufficient condition for the $\mathsf{NP}$-hard problem of relative approximation of marginals, this raises the question whether it is still possible to efficiently approximate the distribution $P$ with a compact PC $Q$ that is tractable for marginals. Unfortunately, we next answer this in the negative.
3.2 Hardness of Approximating Distributions using Tractable Models for Marginals
We consider approximating potentially unnormalized distributions, which can be considered a generalization of probability distributions by omission of the normalizing constant.
Definition 3.3 (Unnormalized Distributions).
Any (unnormalized) distribution $\hat{P}:\operatorname{val}(\mathbf{X})\to\mathbb{R}$ must satisfy the following: (1) $\hat{P}(\mathbf{x})\geq 0$ for any $\mathbf{x}$; (2) The normalization constant $Z=\sum_{\mathbf{x}\in\mathbf{X}}\hat{P}(\mathbf{x})$ is well-defined and finite.
See from this definition that an unnormalized distribution can easily be converted to a probability distribution $P$ if $Z$ is computable in polynomial time: $P(\mathbf{x})=\hat{P}(\mathbf{x})/Z$. Unnormalized distributions are relevant to the goal of tractable approximation, as many probabilistic models that we may want to approximate—including factor graphs [32] and energy-based models [44]—represent unnormalized distributions.